# Spin-Dependent Force and Inverted Harmonic Potential for Rapid Creation of Macroscopic Quantum Superpositions

### v1

@@ VI Fluctuation in magnetic field and wave packet contrast
Since we are considering the effect of classical position and momentum deviations on the contrast, the values of the parameters are the same in $\psi_{\text{L}}(x,t)$ and $\psi_{\text{R}}(x,t)$ except for $x_{c}(t)$ (classical position) and $b(t)$ (associated with classical momentum). The integral of Eq.(37) can be written as:
$C(t)=\exp[-\frac{\Delta x^{2}}{8\sigma_{x}^{2}(t)}-\frac{\sigma_{x}^{2}(t)\Delta b^{2}}{2}],$ (38)
@@
The $x_{\text{L}}$ and $x_{\text{R}}$ are the classical positions of the wave packets of the left and right arms of the interferometer. The $b_{\text{L}}$ and $b_{\text{R}}$ correspond to the values of the parameter $b$ in the left and right wave packets. For Eq.(38), only the exponential decay terms associated with the classical position and momentum deviations are retained while the amplitude and phase factor is neglected [78]. Since both the classical trajectory $x(\eta,t)$ and the parameter $b(\eta,t)$ are functions of the gradient $\eta$ [^7] For a linear magnetic field, $\eta$ is taken as $\eta_{l}$. For a nonlinear magnetic field, $\eta$ is taken as $\eta_{n}$. and time $t$, when there is a small fluctuation in the gradient, Eq.(VI) can be written as:
@@
Note that the change in the parameter $b$ is caused by a change in the classical momentum, so the parameter $b$ needs to be written as an expression with respect to the momentum $p$. Substituting the expressions for $x(\eta,t)$ and $b(\eta,t)$ into Equation (VI) for the HP and IHP cases, and then combining the results with Equation (38), would reveal the effect of fluctuations in the gradient on the contrast in both cases.
VI.1 Wave packet contrast in the case of HP
The expression for the classical momentum in the HP case can be obtained by deriving Eq.(34) with respect to time:
$p(t)=-mx_{0}\omega\sin(\omega t)+p_{0}\cos(\omega t).$ (41)
Considering the Gaussian wave packet as an initial condition (see Eq.(32)) and setting the initial momentum $p_{0}=0$, in combination with Eq.(41), the parameter $b$ (see Eq.(A.1) in the Appendix A) can be written as:
$b(t)=\frac{\hbar p(t)}{4m^{2}\omega^{2}\sigma_{0}^{4}\cos^{2}(\omega t)+\hbar^{2}\sin^{2}(\omega t)}.$ (42)
Combining Eq.(34), (38), (VI), (41) and (42) gives the variation of wave packet contrast with time when there is a small fluctuation $\Delta\eta_{l}$ in the linear magnetic field gradient, as shown in Fig.5. The evolution time is set to 0.2 s to see the changing trend of the contrast more clearly. Over a short period of time, the contrast is generally oscillating and decaying with time. The larger the gradient fluctuation, the faster the decay. The initial wave packet spatial width also has an effect on the contrast. The wider the initial spatial width, the slower the contrast decays.
VI.2 Wave packet contrast in the case of IHP
Figure 6: Wave packet contrast over time when there is a small fluctuation in the nonlinear magnetic field gradient. The red, blue, and orange solid (dashed) lines represent gradient fluctuations of $\Delta\eta_{n}/\eta_{n}=1\times 10^{-4},3\times 10^{-4}\,\,\text{and}\,\,5\times 10^{-4}$, respectively. The solid lines correspond to an initial wave packet spatial width of $\sigma_{0}=2\times 10^{-11}$ m, and the dashed lines correspond to an initial wave packet spatial width of $\sigma_{0}=3\times 10^{-11}$ m. We set the mass $m=10^{-15}$ kg and the gradient $\eta_{n}=1\times 10^{6}$ T/$\text{m}^{2}$.
Derivation of Eq.(36) with respect to time yields the momentum for the IHP case:
$p^{\prime}(t)=mx_{0}\omega\sinh(\omega t)+p_{0}\cosh(\omega t).$ (43)
Consider also the Gaussian wave packet as an initial condition (see Eq.(32)) and set the initial momentum $p_{0}=0$. Combined with Eq.(43), the parameter $b$ (see Eq.(A.2) in the Appendix A) can be written as:
$b(t)=-\frac{\hbar p^{\prime}(t)}{4m^{2}\omega^{2}\sigma_{0}^{4}\cosh^{2}(\omega t)+\hbar^{2}\sinh^{2}(\omega t)}.$ (44)
Combining Eq.(36), (38), (VI), (43) and (44) gives the variation of wave packet contrast with time when there is a small fluctuation $\Delta\eta_{n}$ in the nonlinear magnetic field gradient, as shown in Fig.6. The wave packet contrast decays with time. The smaller the fluctuation $\Delta\eta_{n}$, the slower the decay. Also, increasing the initial wave packet spatial width can slow down the wave packet contrast decay.
For both HP and IHP cases, reducing the fluctuation of the magnetic field gradient and increasing the initial wave packet width can suppress the decay of contrast.
VII Conclusion
@@
In this paper, we investigated the use of spin-dependent force and IHP to generate massive large spatial superposition states and construct a full-loop interferometer. The spin-dependent force enables a massive wave packet to achieve a small spatial separation in a short time. Subsequently, the IHP allows this separation to increase rapidly. For a nanodiamond with a mass of $10^{-15}$ kg, we found that a superposition size of 50 $\mu$m can be achieved in 0.1 s. We also analytically calculated the evolution of the wave packet in both the HP and IHP using path integrals. Additionally, we discussed how fluctuations in the magnetic field gradient affect the wave packet contrast, concluding that smaller gradient fluctuations and a wider wave packet width result in a slower decrease in contrast.
@@
Our study focused on a nanodiamond embedded with an NV centre. The nanodiamond possesses mechanical degrees of freedom (CoM motion), internal degrees of freedom (phonons), and rotational degrees of freedom. Currently, both center-of-mass motion and phonons can be cooled to the ground state [49], and phonons are difficult to excite during movement in HP and IHP [79, 80]. The rotational degree of freedom can affect the final wave packet contrast of the interferometer; however, if the NV center is at or near the nanodiamond center, this effect can be mitigated by fine-tuning the magnetic field and timing [81, 82, 64]. Furthermore, the electron spin coherence time of the NV center can be as long as 1 s at low temperatures and under pure nanocrystal conditions [83, 84, 85]. These findings indicate the feasibility of realizing macroscopic quantum states with masses ranging from $10^{-17}$ to $10^{-14}$ kg and superposition sizes on the order of micrometres in the laboratory.
Note added: Following the completion of our work, we identified corresponding independent research by Sougato Bose’s group (on arXiv today).
@@ acknowledgments
We are grateful to Tian Zhou, Ryan Marshman and Sougato Bose for stimulating discussions. R. Z. and Q. X. are supported by China Scholarship Council (CSC) fellowship. A. M. is supported by Sloan and the Gordon and Betty Moore foundations.
Appendix A Wave packet evolution
@@
A.1 Wave packet evolution in a HP
@@
A.2 Wave packet evolution in a IHP

### v2

@@ V.1 Fluctuation in magnetic field and wave packet contrast
Since we focus on the effect of classical position and momentum deviations on contrast, all parameters in $\psi_{\text{L}}(x,t)$ and $\psi_{\text{R}}(x,t)$ remain the same except for $x_{c}(t)$ (classical position) and $b(t)$ (related to classical momentum). The integral in Eq. (40) can be expressed as:
$C(t)=\exp\left[-\frac{\Delta x^{2}}{8\sigma_{x}^{2}(t)}-\frac{\sigma_{x}^{2}(t)\Delta b^{2}}{2}\right],$ (41)
@@
Here, $x_{\text{L}}$ and $x_{\text{R}}$ are the classical positions of the wave packets in the left and right interferometer arms, respectively. Similarly, $b_{\text{L}}$ and $b_{\text{R}}$ represent the values of parameter $b$ for the left and right wave packets. In Eq. (41), only the exponential decay terms associated with classical position and momentum deviations are retained, while amplitude and phase factors are neglected [75].
Since both the classical trajectory $x(\eta,t)$ and parameter $b(\eta,t)$ depend on the gradient $\eta$ [^9] For a linear magnetic field, $\eta$ corresponds to $\eta_{l}$. For a nonlinear magnetic field, $\eta$ corresponds to $\eta_{n}$. and time $t$, a small fluctuation $\delta\eta$ in gradient modifies Eq. (V.1) as:
@@
The parameter $b$ is related to classical momentum variations, so it can be expressed in terms of momentum $p$. By substituting the expressions for $x(\eta,t)$ and $b(\eta,t)$ into Eq. (V.1) for both HP and IHP cases, and then combining the results with Eq. (41), we can determine the effect of gradient fluctuations on contrast in both scenarios.
In our experimental setup, both linear and nonlinear magnetic fields are used. We analyze their gradient fluctuations separately. First, we assume a small fluctuation $\delta\eta_{l}$ in the linear magnetic field gradient while keeping the nonlinear gradient constant. By combining the equations of motion from stage 1 to stage 5, Eqs. (12), (13), (III), (23), and (III), we compute the position and momentum deviations induced by this gradient fluctuation at each stage.
To calculate the contrast, we also need the wave packet width at the final stage. Since analytical expressions exist for wave packet evolution in HP and IHP (Eqs. (36) and (38)), we iteratively substitute initial conditions and propagate the wave packet width through each stage until the fifth stage is reached. This process is solved numerically, and the resulting contrast variation with linear gradient fluctuations is shown in Fig. 5.
[…]
The procedure for evaluating the effect of nonlinear magnetic field gradient fluctuations on contrast is similar. Here, we assume a fluctuation $\delta\eta_{n}$ in the nonlinear magnetic field gradient while keeping the linear gradient constant. Again, by combining the equations of motion from stage 1 to stage 5, we compute the corresponding position and momentum deviations. Substituting these into Eq. (41) and calculating the resulting wave packet width evolution, we obtain the contrast variation as a function of $\delta\eta_{n}$, shown in Fig. 5.
From Fig. 5, we observe that to maintain 99% contrast, the gradient fluctuation must be below $10^{-7}$ for both linear and nonlinear magnetic fields. Interestingly, nonlinear magnetic fields have also been used to generate a macroscopic spatial superposition state in [54], but without employing IHP. In that case, to maintain 99% contrast, the gradient fluctuation must be below $10^{-9}$. This means that incorporating IHP relaxes the constraint on nonlinear gradient fluctuations by two orders of magnitude. This improvement arises because IHP increases the wave packet width, making the system more tolerant to position and momentum deviations caused by gradient fluctuations.
V.2 Deviation in initial position and wave packet contrast
In this section, we examine how initial position deviations affect the wave packet contrast, given by Eq. (41). Ideally, the classical trajectory is perfectly closed, resulting in $\Delta x=0$ and $\Delta b=0$. However, small deviations in the initial position introduce changes in $\Delta x$ and $\Delta b$, altering the final contrast. Here, we define $\Delta x$ and $\Delta b$ as:
$\Delta x$ $=x(x_{0}+\delta x,t)-x(x_{0},t),$
$\Delta b$ $=b(p(x_{0}+\delta x),t)-b(p(x_{0}),t),$ (44)
where $\delta x$ is a small constant representing the deviation from the initial position. To compare the effect of initial position deviation on contrast in different scenarios, we first analyze the case where only the HP stage is considered, without the influence of the IHP stages. We assume that the wave packet completes a full period in the first HP stage and that the two wave packets recombine at the end of this stage. By substituting the equations of motion for the HP case into Eq.(V.2) and combining them with Eq.(41), we obtain the contrast as a function of the initial position deviation:
$C=\exp\left(-\frac{\delta x^{2}}{8\sigma_{0}^{2}}\right).$ (45)
Here, the contrast depends only on the initial position deviation and the initial wave packet width $\sigma_{0}$. The relationship between contrast and $\delta x$ given in Eq.(45) is shown in Fig.6. In the HP-only case, to maintain a contrast greater than 99%, the initial position deviation must be within the width of the initial wave packet, which is approximately $10^{-11}$ m.
Figure 6: Wave packet contrast as a function of initial position deviations. The solid pink line represents the contrast when the wave packets experienced the IHP stages (the contrast of the final stage). The green dashed line represents the contrast that the wave packets did not experience the IHP stages. At this point, only the first stage is considered and the wave packets are allowed to recombine in the first stage. We take the values of mass $m=10^{-15}$ kg and initial wave packet width $\sigma_{0}=2\times 10^{-11}$ m, and other values listed in Table 2.
Next, we examine how initial position deviations influence the final contrast when the IHP stages are included. The calculation follows a similar approach to that used for analyzing the effect of gradient fluctuations on contrast. Using the equations of motion from stages 1 to 5, we determine the position and momentum deviations caused by the initial position deviation. The wave packet width is then computed numerically by incorporating the wave packet evolution equations for both HP and IHP stages. Finally, by substituting the computed position and momentum deviations, as well as the wave packet width, into Eq.(41), we obtain the contrast as a function of initial position deviation, illustrated in Fig.6.
From Fig.6, we observe that when the IHP stages are included, the initial position deviation can be as large as $10^{-9}$ m while still maintaining a contrast above 99%. This suggests that although the IHP stage amplifies the effect of initial position deviations, its impact on increasing the wave packet width is even more significant. As a result, the requirement for initial position accuracy is relaxed by two orders of magnitude while maintaining the same contrast level, compared to the case without the IHP stages.
VI Conclusion
@@
In this paper, we explored the use of spin-dependent forces and IHP to generate large spatial superposition states of massive objects and construct a full-loop interferometer. The spin-dependent force enables a small initial spatial separation of the massive wave packet within a short time, while the IHP subsequently amplifies this separation significantly. For a nanodiamond with a mass of $10^{-15}$ kg, our analysis shows that a superposition size of approximately 50 $\mu$m can be achieved within 0.1 s. We also provided an analytical treatment of wave packet evolution in both the HP and IHP using path integral methods.
Based on the equations of motion for each stage and the exact wave packet evolution, we analyzed the impact of gradient fluctuations and initial position deviations on the interference contrast. Our results indicate that incorporating the IHP significantly relaxes the experimental constraints. Specifically, the tolerance for nonlinear magnetic field gradient fluctuations improves by two orders of magnitude, from $10^{-9}$ to $10^{-7}$, while maintaining 99% contrast. Similarly, the requirement for position control accuracy is reduced by two orders of magnitude, from $10^{-11}$ m to $10^{-9}$ m, when considering both the HP and IHP stages. This relaxation arises because the IHP enhances wave packet expansion more effectively than it amplifies gradient fluctuations and position deviations, thereby increasing robustness against these imperfections.
@@
Our study focused on a nanodiamond embedded with an NV centre. The nanodiamond possesses mechanical degrees of freedom (CoM motion), internal degrees of freedom (phonons), and rotational degrees of freedom. Currently, the CoM motion can be cooled to the ground state [41], and the internal phonons are difficult to excite during movement in HP and IHP [76, 77]. The rotational degree of freedom can affect the final wave packet contrast of the interferometer; however, if the NV center is at or near the nanodiamond center, this effect can be mitigated by fine-tuning the magnetic field and timing [65, 78, 62]. Furthermore, the electron spin coherence time of the NV center can be as long as 1 s at low temperatures and under pure nanocrystal conditions [79, 80, 47]. These findings indicate the feasibility of realizing macroscopic quantum states with masses ranging from $10^{-17}$ to $10^{-14}$ kg and superposition sizes on the order of micrometres in the laboratory.
Note added: Recent related independent work [81] has come to our attention.
@@ acknowledgments
We are grateful to Tian Zhou, Ryan Marshman, Ryan Rizaldy and Sougato Bose for stimulating discussions. R. Z. and Q. X. are supported by China Scholarship Council (CSC) fellowship. AM’s research is funded in part by the Gordon and Betty Moore Foundation through Grant GBMF12328, DOI 10.37807/GBMF12328. This material is based upon work supported by Alfred P. Sloan Foundation under Grant No. G-2023-21130.
Appendix A One dimensional motion approximation
For a static magnetic field, Maxwell’s equations require that $\divergence{\mathbf{B}}=0$ and ×$\curl{\mathbf{B}}=0$. In our scheme, the complete linear magnetic field is written as
$\mathbf{B}=(B_{0}+\eta_{l}x)\mathbf{e}_{x}-\eta_{l}y\mathbf{e}_{y},$ (46)
where $B_{0}$ is the bias field along the $x$-direction, $\eta_{l}$ is the gradient of the linear field (identical in the $x$ and $y$ directions), and $\mathbf{e}_{x}$, $\mathbf{e}_{y}$ are the unit vectors in the $x$ and $y$ directions, respectively. Similarly, the complete nonlinear magnetic field is given by
$\mathbf{B}=(B_{0}-\eta_{n}x^{2}+\eta_{n}y^{2})\mathbf{e}_{x}-2\eta_{n}xy\mathbf{e}_{y},$ (47)
where $\eta_{n}$ is the nonlinear gradient with unit T/m2. Other symbols have the same meanings as in the linear case. The potential energy for a nanodiamond embedded with an NV center subjected to an external magnetic field is
$U=-\frac{\chi_{\rho}m}{2\mu_{0}}\mathbf{B}^{2}+\hbar\gamma_{e}\mathbf{S}\dotproduct\mathbf{B},$ (48)
where $\chi_{\rho}$ is the mass susceptibility of the diamond, $m$ is its mass, $\mu_{0}$ is the vacuum permeability, $\hbar$ is the reduced Planck constant, $\gamma_{e}$ is the electron gyromagnetic ratio, and $\mathbf{S}$ is the spin of the NV center. By substituting Eqs. (46) and (47) into Eq. (48), we obtain the potential energy corresponding to both linear and nonlinear magnetic fields, as illustrated in Fig. 7.
Force in the Linear Magnetic Field
Substituting Eq. (46) into Eq. (48), the force in the linear magnetic field is
$\mathbf{F}$ $=-\gradient U_{l},$
$=\quantity(\frac{\chi_{\rho}m}{\mu_{0}}\eta_{l}^{2}x+\frac{\chi_{\rho}m}{\mu_{0}}B_{0}\eta_{l}-\hbar\gamma_{e}S_{x}\eta_{l})\mathbf{e}_{x}+\quantity(\frac{\chi_{\rho}m}{\mu_{0}}\eta_{l}^{2}y+\hbar\gamma_{e}S_{y}\eta_{l})\mathbf{e}_{y},$ (49)
where $U_{l}$ is the potential energy associated with the linear field. Table 3 lists typical values of the physical constants. Advances in current technology have enabled the cooling of nanoparticles with masses around $10^{-18}$ kg to their motional ground states [82, 41], and by further improving the environmental conditions, even larger masses (e.g. $10^{-15}$ kg) can be cooled [83]. Notice that in Eq. (A) the force has a $y$-component. However, because a bias magnetic field $B_{0}$ is applied along $x$, the electron spin undergoes Larmor precession about the $x$-axis with frequency $\omega_{L}=\absolutevalue{\gamma_{e}B_{0}}\approx 10^{12}$ Hz, which is much faster than the experimental frequency (approximately 10 Hz). Hence, the average value of $S_{y}$ is zero. Furthermore, the restoring force from the $y$-component creates a simple harmonic motion with an amplitude of about $10^{-11}$ m (the width of the ground state wave packet), which is negligible compared to the superposition size we aim to achieve. In addition, we can further restrict the $y$ motion by applying an external trapping potential. For instance, Ref. [84] employs a three-dimensional magnetic field generated by a current-carrying wire that levitates nanodiamonds in the $z$ direction, confines motion in $y$, and allows free evolution in $x$. Therefore, it is reasonable to consider the nanoparticle motion as one-dimensional (along $x$).
[…]
$\approx 5\times 10^{-18}$ 10 $\approx 1\times 10^{-34}$ $\approx 2\times 10^{11}$ $\approx 5\times 10^{3}$ $\approx 1\times 10^{6}$
Table 3: Values of the physical parameters used for force estimation. Here $B_{0}$ is the bias magnetic field, $\eta_{l}$ and $\eta_{n}$ are the gradients of the linear and nonlinear magnetic fields, respectively. The mass $m$ is taken as $10^{-15}$ kg. $\chi_{\rho}$, $\mu_{0}$, $\hbar$, and $\gamma_{e}$ are fundamental constants.
Figure 7: Comparison of the potential landscapes for the linear and nonlinear magnetic fields. (a) Three-dimensional plot of the potential corresponding to the linear magnetic field. (b) Potential profile in the $x$-direction for the linear magnetic field when $y=0$. The solid pink line corresponds to $S_{x}=1$, while the dashed pink line corresponds to $S_{x}=-1$. (c) Potential profile in the $y$-direction for the linear magnetic field when $x=0$. Since $x=0$, the potentials for spin states $S_{x}=\pm 1$ coincide in the $y$-direction. (d) Three-dimensional plot of the potential corresponding to the nonlinear magnetic field. (e) Potential profile in the $x$-direction for the nonlinear magnetic field when $y=0$. (f) Potential profile in the $y$-direction for the nonlinear magnetic field when $x=0$. For visualization purposes, we set $B_{0}=0$ in the linear magnetic field case. In the nonlinear case, we introduce a reference potential $U_{0}=1539.84$ eV and define $\Delta U=U-U_{0}$, ensuring that the potential variation starts from zero, as shown in (e) and (f). The values of other physical parameters are listed in Table 3.
Force in the Nonlinear Magnetic Field
By substituting Eq. (47) into Eq. (48), the force in the nonlinear magnetic field is given by
$\mathbf{F}$ $=-\gradient U_{nl},$
$=\quantity(\frac{2\chi_{\rho}m}{\mu_{0}}\eta_{n}^{2}x^{3}-\frac{2\chi_{\rho}m}{\mu_{0}}B_{0}\eta_{n}x+2\hbar\gamma_{e}S_{x}\eta_{n}x+\frac{2\chi_{\rho}m}{\mu_{0}}\eta_{n}^{2}y^{2}x+2\hbar\gamma_{e}S_{y}\eta_{n}y)\mathbf{e}_{x}$
$\quad+\quantity(\frac{2\chi_{\rho}m}{\mu_{0}}\eta_{n}^{2}y^{3}+\frac{2\chi_{\rho}m}{\mu_{0}}B_{0}\eta_{n}y-2\hbar\gamma_{e}S_{x}\eta_{n}y+\frac{2\chi_{\rho}m}{\mu_{0}}\eta_{n}^{2}x^{2}y+2\hbar\gamma_{e}S_{y}\eta_{n}y)\mathbf{e}_{y},$ (50)
where $U_{nl}$ is the potential associated with the nonlinear field. During the nonlinear field stage, the spin states transform to |$\ket{S_{x}}=\ket{0}$ and |$\ket{S_{y}}=(\ket{+1}+\ket{-1})/\sqrt{2}$. Due to the bias magnetic field $B_{0}$ along $x$, the $S_{y}$ component undergoes rapid Larmor precession, leading to an average value of zero. Consequently, the terms involving $S_{x}$ and $S_{y}$ in Eq. (A) can be neglected. Substituting parameter values from Table 3 into Eq. (A) [^10] Here, $x$ is taken up to 25 $\mu$m, considering that the maximum superposition size is around 50 $\mu$m., we find that the dominant term in the $x$-direction is $(-2\chi_{\rho}m/\mu_{0})B_{0}\eta_{n}x$. This term corresponds to the IHP, as shown in Fig. 7(e), and is responsible for the rapid separation of the wave packets. Similarly, the dominant term in the $y$-direction is $(-2\chi_{\rho}m/\mu_{0})B_{0}\eta_{n}y$,which corresponds to HP, as illustrated in Fig. 7(f). This term effectively restricts motion in the $y$-direction. If the initial state is the ground state, then the motion of the wave packet in the $y$-direction follows simple harmonic oscillation with an amplitude comparable to the width of the initial wave packet (approximately $10^{-11}$ m). Since this amplitude is negligible in comparison to the superposition size, the motion along $y$ can be effectively disregarded. As a result, during the nonlinear field stage, the nanoparticle’s motion remains predominantly confined to the $x$-direction. The effective potential in this regime ensures rapid separation of the wave packets along $x$, while simultaneously suppressing displacement along $y$.
In summary, both for linear and nonlinear magnetic field configurations, the nanoparticle experiences forces with components in both $x$ and $y$. However, due to the rapid Larmor precession along $x$ (causing the mean $S_{y}$ to vanish) and an effective trapping potential that confines $y$-motion, we can safely approximate the nanoparticle motion as one-dimensional along the $x$-axis.
Appendix B Wave packet evolution
@@
B.1 Wave packet evolution in a HP
@@
B.2 Wave packet evolution in a IHP